# Enhanced Athermal Phonon Responsivity in a Kinetic Inductance Detector with Integrated Phonon Collectors

### v1

@@ Abstract
The simulations were calibrated against two identical resonators but in Aluminum. We obtained $\alpha=18.9\,\%$ and $\alpha=15.6\,\%$ for the FunKID and the KID, respectively, much larger than those obtained in Aluminum, $\alpha=7.0\,\%$ and $\alpha=5.0\,\%$ using the same design, consistently with an increase of kinetic inductance $L_{k}$ Cardani et al. (2018).
We also measured the feed-line critical temperature $T_{c}=(0.83\pm 0.08)\,\mathrm{K}$, corresponding to a superconducting gap $\Delta_{0}=(125.9\pm 12.6)\,\mathrm{\mu eV}$, and we assumed that this coincided with the critical temperature of the normal KID. Such value is in reasonable agreement with that found by the CALDER project Cardani et al. (2018) of $0.81\,\mathrm{K}$, using a similar film. Nevertheless such estimation is reliable as long as it is applied to the normal KID. The presence of funnels is expected to modify the FunKID critical temperature, but in the current setup we do not have a direct way to measure it.
We also measured $\delta f_{r}/f_{r}$, the relative shift of the resonant frequency with respect to the base temperature resonant frequency, $f_{r}$. Using a model based on the BCS theory Gao et al. (2008), it is possible to write:
$\frac{\delta f_{r}}{f_{r}}=-\frac{\alpha}{2}S_{2}(f_{r},T)\sqrt{\frac{{2\pi k_{B}T}}{\Delta(T)}}e^{-\Delta(T)/k_{B}T}$ (6)
where $\Delta(T)$ is the AlTiAl superconducting order parameter Gao et al. (2008). For this purpose, we measured the resonant frequency of each resonator at different temperatures, while keeping the transmission bias powers fixed at $87\,\mathrm{dBm}$ and $71\,\mathrm{dBm}$ for the FunKID and the KID, respectively. These powers are lower than the operative calibration powers in order to isolate the effect of thermally generated excess QPs in the resonators, while keeping the micro-wave induced contribution as low as possible de Visser et al. (2012). However, potential limitations may arise when applying such a model to inhomogeneous and multilayer superconductors. In order to test the applicability of such model to our devices, we compared the data with the BCS model calculated with our best predictions of $\alpha$ and $\Delta_{0}$ (Fig. 5).
@@
To calibrate the detectors, the FunKID and the KID were operated at bias transmission calibration powers $P_{\mathrm{{cal}}}$ of $68\,\mathrm{dBm}$ and $62\,\mathrm{dBm}$, respectively, at which they exhibited good performance in terms of signal to noise ratio Cardani et al. (2018). The choice to operate the FunKID with a lower calibration power was motivated by the observation that on the FunKID non-linear effects became relevant at smaller power than the KID Swenson et al. (2013). Optical pulses of increasing total energy from $\sim 9\,\mathrm{keV}$ up to $\sim 90\,\mathrm{keV}$ were fired onto the back of the substrate by means of the central fiber to characterize the response of the two KIDs Del Castello (2024). Finally, the signals were processed offline with a matched filter Gatti and Manfredi (1986) in order to further improve the signal to noise ratio.
Two pulses generated by an energy release of around $90\,\mathrm{keV}$ into the substrate have been superimposed in Fig. 6. It is possible to observe that the FunKID exhibits a pulse response about 5.5 times higher than the KID. Our best prediction for the gain in responsivity under identical electrical conditions was around 8. According to Eq. 4 the dominant degradation of the FunKID response might be associated with its smaller quality factor $Q$, which is a factor 1.9 lower than that of the KID. However, our best prediction of $\alpha$ of the FunKID is about 1.2 higher than that of the KID. By correcting the best gain prediction with these two effects, we obtain a gain of around 5.1, similar to measurements.
@@
The decay times $\tau_{d}$ of the FunKID and KID are about $0.15\,\mathrm{ms}$ and $0.45\,\mathrm{ms}$, respectively. The longer decay time of the KID is in line with the higher internal quality factor Cruciani et al. (2016). The rise times $\tau_{\mathrm{rise}}$ amount to around $12\,\mathrm{\mu s}$ and $20\,\mathrm{\mu s}$, respectively. These are significantly higher than the respective ring times (see Tab 1). In addition, the rise time of the FunKID is higher than the QPs diffusion time scale $\tau_{\mathrm{diff}}\simeq 4\,\mathrm{\mu s}$ previously defined. It seems reasonable that in both cases the rise times of the pulses are dominated by the phonon absorption time $\tau_{\mathrm{ph}}$ Swenson et al. (2010); Martinez et al. (2019), which can be faster in the case of the FunKID due to its higher phonon collection efficiency.
@@
Figure 6: Two pulses generated by the same energy release. A higher response is observed for the FunKID resonator compared to the normal KID. Also, it is possible to notice the higher $\tau_{d}$ of the KID pulse, which takes almost $2\,\mathrm{ms}$ to come back to the baseline level.
@@
The results of the energy calibrations Cardani et al. (2018); Delicato et al. (2025); Del Castello (2024) of the two KIDs, the phase responsivity $r$, the noise RMS after applying the matched filter, $\sigma_{N}$, and the corresponding baseline energy resolution of the device, $\sigma_{E}$, have been reported in Tab. 2. The phase responsivity is about 5.3 higher at the calibration power for the FunKID, which is in line with the measured responsivity gain previously discussed. Moreover, an improvement of around a factor of 2 for the energy resolution in $\mathrm{eV}$ is measured, despite the worse noise RMS $\sigma_{N}$ of the FunKID. The improvement of the FunKID responsivity was therefore proven, even with a total quality factor smaller by a factor of about 2.
@@
While the tested device served as a proof of concept and the improvement in detector responsivity was demonstrated, an enhancement in energy resolution is still required, as the current performance remains worse than previously achieved results Delicato et al. (2024). Achieving this improvement will demand further optimization, such as an additional increase of the total quality factors $Q$, which in turn requires higher internal quality factors $Q_{i}$.

### v2

@@ Abstract
The simulations were calibrated against two identical resonators but in Aluminum. We obtained $\alpha=18.9\,\%$ and $\alpha=15.6\,\%$ for the FunKID and the KID, respectively, much larger than those obtained in Aluminum, $\alpha=7.0\,\%$ and $\alpha=5.0\,\%$ using the same design.
We measured a feed-line $T_{c}=(0.83\pm 0.08)\,\mathrm{K}$, corresponding to $\Delta_{0}=(125.9\pm 12.6)\,\mathrm{\mu eV}$, in reasonable agreement with the $0.81\,\mathrm{K}$ reported in Cardani et al. (2018). We assumed that this value applies to both the resonators, although funnels likely modify the FunKID critical temperature, which cannot be measured directly in the presented setup.
We measured the resonant frequency shift in temperature relative to the base temperature resonant frequency, $\delta f_{r}/f_{r}$, with bias powers fixed at $87\,\mathrm{dBm}$ (FunKID) and $71\,\mathrm{dBm}$ (KID), chosen to minimize the microwave-induced QPs excess de Visser et al. (2012). We compared the data to a model based on the BCS theory Gao et al. (2008), calculated with the predicted $\alpha$ and $\Delta_{0}$ previously discussed (Fig. 5). However, since the BCS theory might be limited in case of inhomogeneous and multilayer superconductors, the comparison is qualitative and aims to test the applicability of the model.
@@
To calibrate the detectors, the FunKID and the KID were operated at bias transmission calibration powers $P_{\mathrm{{cal}}}$ of -68 dBm and -62 dBm, respectively, at which they exhibited good performance in terms of signal to noise ratio Cardani et al. (2018). Optical pulses of increasing total energy from around 9 keV up to 90 keV were fired onto the back of the substrate by means of the central fiber Del Castello (2024). Finally, signals were processed offline with a matched filter Gatti and Manfredi (1986) in order to further improve the signal to noise ratio.
@@
Two pulses generated by an energy release of around 90 keV into the substrate have been superimposed in Fig. 6. The FunKID exhibits a pulse response about 5.5 times larger than the standard KID. This value is lower than the prediction of 8 derived assuming identical electrical conditions. However, by correcting this estimate using Eq. 4 with the measured ratios of quality factors and kinetic inductance fractions (approximately 0.53 and 1.21, respectively), we obtain a gain of around 5.1, similar to measurements. The decay times $\tau_{d}$ of the FunKID and KID are about 0.15 ms and 0.45  ms, respectively. The smaller decay time of the FunKID is in line with the smaller internal quality factor Cruciani et al. (2016). The rise times $\tau_{\mathrm{rise}}$ amount to around 12 $\mathrm{\mu s}$ and 20 $\mathrm{\mu s}$, respectively. These are significantly higher than the respective ring times (see Tab 1). In addition, the rise time of the FunKID is higher than the QPs diffusion time scale $\tau_{\mathrm{diff}}\simeq 4\,\mathrm{\mu s}$ previously defined. It seems reasonable that in both cases the rise times of the pulses are dominated by the phonon absorption time $\tau_{\mathrm{ph}}$ Swenson et al. (2010); Martinez et al. (2019), which can be faster in the case of the FunKID due to its higher phonon collection efficiency.
@@
Figure 6: Two pulses generated by the same energy release. A higher response is observed for the FunKID resonator compared to the normal KID.
@@
The results of the energy calibrations Cardani et al. (2018); Delicato et al. (2025); Del Castello (2024) have been reported in Tab. 2. The responsivity is about 5.3 higher at the calibration power for the FunKID, which is in line with the measured responsivity gain previously discussed. Moreover, an improvement of around a factor of 2 for the energy resolution in eV is measured, despite the worse noise RMS $\sigma_{N}$ of the FunKID. The improvement of the FunKID responsivity was therefore proven, even with a smaller total quality factor.
@@
While the tested device served as a proof of concept and the improvement in detector responsivity was demonstrated, an enhancement in energy resolution is still required, as the current performance remains worse than previously achieved results Delicato et al. (2024). Achieving this improvement demands further optimization, such as an additional increase of the total quality factors $Q$, which in turn requires higher internal quality factors $Q_{i}$.